the following information was obtained from a package of lean hotdogs. one hot dog: fat equals 7 g. protein equals 12 G. carbohydrates equals 0 grams. from this information find a percentage of calories from fat in this hot dog
Answer #1

Carbohydrates provide 4 calorie/gram

Protein provides 4 calorie/gram

Fat provides 9 calories/gram

Fats = 7 * 9 = 63

Protein = 12* 4= 48

Carbohydrates = 0 * 4 = 0

Total = 63 + 48 = 111 calories

% calorie from fat = (63 / 111) * 100 = 56.75 %